OpenStudy (anonymous):

please solve:

OpenStudy (anonymous):

\[-46\pm \sqrt{46^{2}-4(4)(-130)} \over 2(4)\]

OpenStudy (lgbasallote):

quadratic formula?

OpenStudy (anonymous):

yup :-)

OpenStudy (anonymous):

please solve? er, just type it into a calculator?

OpenStudy (lgbasallote):

no calculator? lol

OpenStudy (anonymous):

need to show steps :(

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

hey .. just solve!

OpenStudy (lgbasallote):

lol what the heck are you supposed to show there =))

hartnn (hartnn):

2.34 and -13.84

OpenStudy (lgbasallote):

you're gonna square 46 manually? =))

OpenStudy (anonymous):

well, first simplify the part inside the square root

OpenStudy (anonymous):

-46 + - 64.77

OpenStudy (lgbasallote):

this thing is too big to actually show steps :))

OpenStudy (anonymous):

first find the sqrt of the number

hartnn (hartnn):

more accurately:2.34706 and -13.84706

OpenStudy (anonymous):

okok, thanks :)

OpenStudy (anonymous):

wait, do you need exact form? (sqrt(2) type of thing)

OpenStudy (anonymous):

yup, if u can :)

OpenStudy (anonymous):

so you can use a calculator too right, (or can you square 46 mentally)

OpenStudy (lgbasallote):

\[\sqrt{(46)^2 - 4(4)(-130)} \implies \sqrt{2,116 + 2080} \implies \sqrt{4,196}\] lol

hartnn (hartnn):

-46+-(sqrt(4196))/8 :P

OpenStudy (lgbasallote):

i still cant believe you have to show this step by step =))

OpenStudy (lgbasallote):

you can simplify it gurther @hartnn \[\implies \frac{-46 \pm 2\sqrt{1049}}{8} \implies \frac{-23 \pm \sqrt{1049}}{4}\]

OpenStudy (lgbasallote):

further*

OpenStudy (anonymous):

=-=''' i don't want to also TxT

OpenStudy (lgbasallote):

well i think that's the simplest you can get without going exact..

OpenStudy (anonymous):

um, isn't that exact already?

OpenStudy (lgbasallote):

by exact i mean the answers you said earlier

OpenStudy (lgbasallote):

with decimal thingies

OpenStudy (anonymous):

approximate

OpenStudy (anonymous):

thanks, guys :)

OpenStudy (anonymous):

Sure, if you're done, can you please close the thread. Thanks

OpenStudy (anonymous):

okok